OpenStudy (anonymous):

the ordered pair (0,0)(1,4)(2,16)(3,36)(4,64). write a rule that represents the function.

OpenStudy (anonymous):

what I had made sense until I reached the part where it said (3,36) and (4,64)

OpenStudy (anonymous):

What do you notice about the y-coordinates for each ordered pair?

OpenStudy (anonymous):

it multiplys by 2s

OpenStudy (math&ing001):

I'm guessng the 4th one is not correct. Should be (3,32)

OpenStudy (anonymous):

yeah i got confused by that one

OpenStudy (anonymous):

No. Let's look at the y-coordinates: 0, 4, 16, 36, 64 Anything interesting about these numbers?

OpenStudy (anonymous):

multiples of 4

OpenStudy (math&ing001):

Sry my bad

OpenStudy (anonymous):

I got confused because of the answer I had before

OpenStudy (math&ing001):

So what do you think the function should be hannah?

OpenStudy (anonymous):

I htought it was 4x^2

OpenStudy (anonymous):

thought*

OpenStudy (anonymous):

Are they multiple of 4 in any order? No. Multiples of 4: {4,8,12,16,20,24,28,...} So those numbers are a subset (part of) the set oif multiples of 4. Try again...what is common to all those y-coordinate numbers ?

OpenStudy (anonymous):

want a hint?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

They are all perfect squares. But the question is, what pattern do these perfect squares form?

OpenStudy (anonymous):

2,4,6,and 8

OpenStudy (anonymous):

I assume mean, 0^2 = 0 2^2 = 4 4^2 = 16 6^2 = 36 8^2 = 64 So what exactly is the pattern???

OpenStudy (anonymous):

Notice the y-coordinates are 0,4,16,36,64, ....

OpenStudy (anonymous):

they all have an exponent of 2?

OpenStudy (anonymous):

well, yes, that's what we mean by a perfect square. To see if you understand the pattern, what will be the coordinates of the point AFTER (4,64)?

OpenStudy (anonymous):

(5,81)?

OpenStudy (anonymous):

No. Try again.

OpenStudy (anonymous):

sadly idk T_T

OpenStudy (math&ing001):

try getting y-coordinate as function of the x-coordinate

OpenStudy (anonymous):

0^2 = 0 ;2^2 = 4 ; 4^2 = 16; 6^2 = 36; 8^2 = 64 If you notice that it skips by 2, 0^2, then 4^2, then 6^2, then 8^2, so I expect the the next coordinate would be (5,100)...Would you agree?

OpenStudy (anonymous):

Oh now I understand

OpenStudy (anonymous):

First y - coordinate is: 0^2 Second y-coordinate is: 2^2 Third y-coordinate is: 4^2 Fourth y-coordinate is: 6^2 Fifth y-coordinate is: 8^2 Sixth y-coordinate is: ???????

OpenStudy (anonymous):

Sith y-coordinate is: 10^2 So what would be the Seventh y-coordinate?

OpenStudy (anonymous):

12^2

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

Now, let's try to get a rule that generates the coordinates. That was your original question. Can you give me the rule?

OpenStudy (anonymous):

the the rule to represent the function

OpenStudy (anonymous):

Yes, we are looking for a rule that relates y with x. Look at the x-coordinates and the y-coordinates. How are they related?

OpenStudy (anonymous):

I can see how they are related separatly but not together.

OpenStudy (anonymous):

Try looking at it again for a moment or two...see if you can see how they are related together.........

OpenStudy (anonymous):

1*4=4 2*8=16 3*12=36 4*16=64

OpenStudy (anonymous):

very interesting.....

OpenStudy (anonymous):

so how would you write your rule, like y = something with x

OpenStudy (math&ing001):

4*1^2=4 4*2^2=16 4*3^2=36 now just express it with x instead of numbers

OpenStudy (anonymous):

I was looking at the rule from a different angle..with perfect squares

OpenStudy (anonymous):

@HannahKimiko So what would be the rule?

OpenStudy (anonymous):

4x^2?

OpenStudy (math&ing001):

exactly

OpenStudy (math&ing001):

Now what do you think the pair should be ?

OpenStudy (anonymous):

Nice job, @HannahKimiko So what is the rule? Let's write it mathematically as an ordered pair (x,y)?

OpenStudy (anonymous):

y=4x^2?

OpenStudy (anonymous):

Yes, thats perfect...as the question just asks for the rule as a function, and y = 4x^2 is a function of x.

OpenStudy (anonymous):

Yay!!!!!!!!

OpenStudy (anonymous):

And, by the way, since we have the rule...we can easily figure out any coordinate...say when x = 20, then y = 4(20^2) = 4(400) = 1,600.

OpenStudy (anonymous):

thats's interesting ^^

OpenStudy (anonymous):

@HannahKimiko Now, whenever you see such a question, try to see the pattern, and relate that pattern with x, to develop a function rule.

OpenStudy (anonymous):

how do you do one with fractions though?

OpenStudy (anonymous):

Sometimes you do find questions...where they give you a certain pattern, and will ask, "What would be the 10th term, etc..". That's easy once you have the rule.

OpenStudy (anonymous):

Well, the same way with fractions...if you have an exercise like this one that contains fractions, we can look at that if you want.

OpenStudy (anonymous):

um... (1,2/3) (2,4/9)(3,8/27)(4,16/81)(5,32/243)

OpenStudy (anonymous):

OK..take a few moments and see if you can first establish a pattern.

OpenStudy (anonymous):

can you change it to a decimal?

OpenStudy (anonymous):

Dont change into a decimal...look at these fractions 2/3, 4/9, 8/27, 16/81, ... anything seem interesting about these numbers?

OpenStudy (anonymous):

the denominator is a multiple of 3

OpenStudy (anonymous):

and the numerator is a multiple of 2

OpenStudy (anonymous):

so look at the first 2 numbers, 2/3 and 4/9...how are they related?

OpenStudy (anonymous):

it's simplified

OpenStudy (anonymous):

Oh yea? When you simplify 4/9, you get 2/3?

OpenStudy (anonymous):

How so?

OpenStudy (anonymous):

oops lol

OpenStudy (anonymous):

Hint: First term, 2/3 = (2/3)^1 Second term, 4/9 = (2/3)^2 Do you see the pattern now?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So what is the function rule that relates y with x?

OpenStudy (anonymous):

y = ?

OpenStudy (anonymous):

y=2/3^2

OpenStudy (anonymous):

I dont see y in terms of x

OpenStudy (anonymous):

y=2/3^x

OpenStudy (anonymous):

First term: y = (2/3)^1 Second term: = = (2/3)^2 Third term: y = (2/3)^3 So the general rule is what? y = what?

OpenStudy (anonymous):

Yes, y = (2/3)^x. Perfect! So you're getting the hang of it. Just practice, and you'll do fine. If you need more help, write back.

OpenStudy (anonymous):

can you give me a hint?

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

i htought i got it wrong

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

You can see that it is correct. How? Take the 4th term, is the fourth term y = (2/3)^4? (2/3)^4 = 16/81..so it checks.

OpenStudy (anonymous):

And you can find any term, once you know the rule.

OpenStudy (anonymous):

@HannahKimiko You're welcome.

OpenStudy (anonymous):

for the ordered paris (1,2)(2,16)(3,54)(4,128)(5,250) the function is y=2x^3, right?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

what do you do when the numbers are increasing and decreasing?

OpenStudy (anonymous):

likr the pattern is (0,0) (1,.5)(2,2)(3,4.5)(4,8)

OpenStudy (anonymous):

You may need X+1 or x-1, etc...

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

can you help me with this one?

OpenStudy (anonymous):

The last one you typed is an interesting one.

OpenStudy (anonymous):

Try writing a rule for that.

OpenStudy (anonymous):

First, study the pattern, and see if you "can see" ther pattern. Once you see the pattern, write the rule.

OpenStudy (anonymous):

i got you add.5 the 1.5 each time after that.....?

OpenStudy (anonymous):

agree...and then you multiply that by x.

OpenStudy (anonymous):

Now, you must relate y with x, and write a rule. And use the rule to check if that rule works for the given coordinates.

OpenStudy (anonymous):

I will leave that for you to do.

OpenStudy (anonymous):

I will leave that for you to do.